# An Inexact Bregman Proximal Difference-of-Convex Algorithm with Two Types of Relative Stopping Criteria

### v1

@@ Abstract
In this paper, we consider a class of difference-of-convex (DC) optimization problems, where the global Lipschitz gradient continuity assumption on the smooth part of the objective function is not required. Such problems are prevalent in many contemporary applications such as compressed sensing, statistical regression, and machine learning, and can be solved by a general Bregman proximal DC algorithm (BPDCA). However, the existing BPDCA is developed based on the stringent requirement that the involved subproblems must be solved exactly, which is often impractical and limits the applicability of the BPDCA. To facilitate the practical implementations and wider applications of the BPDCA, we develop an inexact Bregman proximal difference-of-convex algorithm (iBPDCA) by incorporating two types of relative-type stopping criteria for solving the subproblems. The proposed inexact framework has considerable flexibility to encompass many existing exact and inexact methods, and can accommodate different types of errors that may occur when solving the subproblem. This enables the potential application of our inexact framework across different DC decompositions to facilitate the design of a more efficient DCA scheme in practice. The global subsequential convergence and the global sequential convergence of our iBPDCA are established under suitable conditions including the Kurdyka-Łojasiewicz property. Some numerical experiments on the $\ell_{1-2}$ regularized least squares problem and the constrained $\ell_{1-2}$ sparse optimization problem are conducted to show the superior performance of our iBPDCA in comparison to existing algorithms. These results also empirically verify the necessity of developing different types of stopping criteria to facilitate the efficient computation of the subproblem in each iteration of our iBPDCA.
@@ 1 Introduction
Due to the significance and popularity of the problem (1.1), tremendous efforts have been made to solve it efficiently, particularly when the problem involves a large number of variables. A classical algorithm for solving problem (1.1) is the so-called DC algorithm (DCA) [34], which iteratively replaces the concave part of the objective with a linear majorant approximation and solves the resulting subproblem as follows: $\bm{x}^{k+1}\in\min\limits_{\bm{x}\in{\mathcal{Q}}}\left\{P_{1}(\bm{x})+f(\bm{x})-\langle\bm{\xi}^{k},\,\bm{x}-\bm{x}^{k}\rangle\right\},$ where $\bm{\xi}^{k}\in\partial P_{2}(\bm{x}^{k})$ is a (classical) subgradient of $P_{2}$ at $\bm{x}^{k}\in\mathcal{Q}$ (see the next section for the definition). But solving this subproblem is generally computationally demanding. By further exploring the structures of the objective, a so-called proximal DC algorithm (pDCA) (see, e.g, [14, 35]) can be adopted to solve problem (1.1), which, in each iteration, not only replaces the concave part of the objective with a linear majorant approximation, but also replaces the smooth convex part with a quadratic majorant approximation. The basic iterative step of the pDCA reads as follows:
@@
The success of the pDCA and the pDCAe also motivates many subsequent works focusing on them and their variants for solving problem (1.1) under appropriate settings; see, for example, [22, 25, 26, 28, 33, 36, 42, 43]. Among them, Takahashi et al. [42] considered the Bregman distance $\mathcal{D}_{\phi}$ (defined in the next section) associated with a kernel function $\phi$ as a proximity measure, and generalized the pDCA to a Bregman proximal DC algorithm (BPDCA), whose basic iterative step reads as follows:
@@
where $\bm{\xi}^{k}\in\partial P_{2}\left(\bm{x}^{k}\right)$. The development of the BPDCA is indeed inspired by recent significant advancements in the Bregman proximal gradient method (see, e.g, [5, 10, 24, 45]). In comparison to the pDCA, such an extended Bregman framework presents two notable advantages. First, the BPDCA can be developed based on the notion of $L$-smooth adaptable property [10, Definition 2.2], which is less restrictive than the global Lipschitz gradient continuity of $\nabla f$ that is required by the pDCA(e) and its variants. Therefore, the BPDCA is applicable for a wider range of problems. Second, by selecting an appropriate kernel function $\phi$, the BPDCA can better leverage the inherent geometry and structure of the problem, potentially leading to a more tractable subproblem (1.2); see examples in [42, 43].
@@
Despite the aforementioned advantages, the BPDCA still encounters limitations in practical implementations. Indeed, although the kernel function can possibly capture the inherent geometry and structure of the problem, the associated subproblem in form of (1.2) generally has no closed-form solution and its computation may still be numerically demanding; see examples in [12, 39, 49] and our numerical section 5. Therefore, for the BPDCA to be implementable and practical, it should allow approximate solutions to the subproblem with progressive accuracy and the corresponding stopping criterion should be practically verifiable. However, to our knowledge, research on the inexact version of the BPDCA remains unexplored. On the other hand, allowing the approximate minimization of the subproblem in proximal DC-type algorithms using the classical or weighted quadratic proximal term (corresponding to $\phi(\cdot):=\frac{1}{2}\|\cdot\|^{2}$ or $\phi(\cdot):=\frac{1}{2}\|\cdot\|^{2}_{H}$ with a symmetric positive definite matrix $H$) has been widely investigated under different (typically simpler) contexts in previous works; see, for example, [22, 28, 32, 40, 41, 44, 56, 57]. However, these inexact versions either require the exact computation of an element in $\partial P_{1}$ (e.g., [22, 28, 40, 44, 56]), which could be difficult to satisfy when $P_{1}$ is not a simple function (see an example in the numerical section 5.2), or they lack the global sequential convergence result (e.g., [32, 41, 57]). Moreover, all these inexact versions require the global Lipschitz continuity of $\nabla f$ when they explicitly approximate the smooth part $f$ by a quadratic majorant. The aforementioned inadequacies could potentially restrict the applicability of these existing inexact frameworks, even under the Euclidean setting.
@@
In this paper, to facilitate the practical implementations of the BPDCA and complement existing inexact proximal DC-type algorithms (see, e.g., [22, 28, 32, 40, 41, 44, 56, 57]), we attempt to develop a general inexact Bregman proximal DC algorithm (iBPDCA) for solving the general problem (1.1). As we shall see later in Section 3, our inexact framework is developed based on two types of relative stopping criteria (SC1) and (SC2), which vary in their strategy to control the error incurred in the inexact minimization of the subproblem. Although the difference may look subtle at the first glance, it indeed influences the behavior of our iBPDCA in both theoretical analysis and computational performance; see more discussions following Algorithm 1 as well as Remarks 5.1 and 5.2. Moreover, the resulting inexact Bregman DCA framework is rather broad to cover many existing exact and inexact methods (studied in, e.g., [40, 41, 42, 44, 56]) in addition to introducing new inexact variants. It can also accommodate different types of errors that may occur when solving the subproblem. The inherent versatility and flexibility of our iBPDCA readily enable its potential application across different DC decompositions for a given problem, thereby facilitating the design of a more efficient DCA scheme, as deliberated in Remark 5.1. Furthermore, in Section 4, we meticulously establish both global subsequential and global sequential convergence results for our iBPDCA under suitable conditions including the renown Kurdyka-Łojasiewicz property. Finally, we conduct some numerical experiments in Section 5 to illustrate the performance of our iBPDCA for solving the $\ell_{1-2}$ regularized least squares problem and the constrained $\ell_{1-2}$ sparse optimization problem. The computational results demonstrate the superior performance of our inexact framework in comparison to existing algorithms, and they also empirically verify the necessity of developing different types of stopping criteria for solving the subproblem.
@@ 2 Notation and preliminaries
In this paper, we present scalars, vectors and matrices in lower case letters, bold lower case letters and upper case letters, respectively. We also use $\mathbb{R}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{m\times n}$ to denote the set of real numbers, $n$-dimensional real vectors, and $m\times n$ real matrices, respectively. For a vector $\bm{x}\in\mathbb{R}^{n}$, $x_{i}$ denotes its $i$-th entry, $\|\bm{x}\|$ denotes its Euclidean norm, $\|\bm{x}\|_{1}$ denotes its $\ell_{1}$ norm defined by $\|\bm{x}\|_{1}:=\sum^{n}_{i=1}|x_{i}|$, and $\|\bm{x}\|_{H}:=\sqrt{\langle\bm{x},\,H\bm{x}\rangle}$ denotes its weighted norm associated with a symmetric positive definite matrix $H$. For a closed set $\mathcal{X}\subseteq\mathbb{E}$, its indicator function $\iota_{\mathcal{X}}$ is defined by $\iota_{\mathcal{X}}(\bm{x})=0$ if $\bm{x}\in\mathcal{X}$ and $\iota_{\mathcal{X}}(\bm{x})=+\infty$ otherwise. The distance from a point $\bm{x}$ to $\mathcal{X}$ is defined by $\mathrm{dist}(\bm{x},\,\mathcal{X}):=\inf_{\bm{y}\in\mathcal{X}}\|\bm{y}-\bm{x}\|$.
@@
When $h$ is continuously differentiable or convex, the above subdifferential coincides with the classical concept of derivative or convex subdifferential of $h$; see, for example, [38, Exercise 8.8] and [38, Proposition 8.12]. Moreover, if $h$ has several groups of variables, we use $\partial_{\bm{x}_{i}}h$ (resp., $\nabla_{\bm{x}_{i}}h$) to denote the partial subdifferential (resp., gradient) of $h$ with respect to the group of variables $\bm{x}_{i}$. In addition, for a proper closed function $h:\mathbb{E}\rightarrow(-\infty,\infty]$ and $\nu>0$, the proximal mapping of $\nu h$ at $\bm{y}\in\mathbb{E}$ is defined by
@@
For a proper closed convex function $h:\mathbb{E}\rightarrow(-\infty,\infty]$ and a given $\varepsilon\geq 0$, the $\varepsilon$-subdifferential of $h$ at $\bm{x}\in{\rm dom}\,h$ is defined by $\partial_{\varepsilon}h(\bm{x}):=\{\bm{d}\in\mathbb{E}:h(\bm{y})\geq h(\bm{x})+\langle\bm{d},\,\bm{y}-\bm{x}\rangle-\varepsilon,~{}\forall\,\bm{y}\in\mathbb{E}\}$, and when $\varepsilon=0$, $\partial_{\varepsilon}h$ simply reduces to the classical convex subdifferential $\partial h$. The function $h$ is essentially smooth if (i) $\mathrm{int}\,\mathrm{dom}\,h$ is non-empty; (ii) $h$ is differentiable on $\mathrm{int}\,\mathrm{dom}\,h$; (iii) $\|\nabla h(x_{k})\|\to\infty$ for every sequence $\{x_{k}\}$ in $\mathrm{int}\,\mathrm{dom}\,h$ converging to a boundary point of $\mathrm{int}\,\mathrm{dom}\,h$; see [37, page 251].
@@
Some remarks are in order concerning this definition. The above restricted $L$-smooth adaptable property is derived by modifying the original $L$-smooth adaptable property introduced in [10, Section 2.2] by imposing a restricted set $\mathcal{X}$, and it readily reduces to the original notion when $\mathcal{X}=\mathbb{E}$. Implementing such a restriction would facilitate the expansion of the original notion to encompass a broader range of choices of $(f,\,\phi)$ with a proper $\mathcal{X}$. For example, when $\nabla f$ is $L_{f}$-Lipschitz continuous on $\mathbb{R}^{n}$ and $\phi$ is $\mu_{\phi}$-strongly convex on $\mathcal{X}\subseteq\mathbb{R}^{n}$, then it can be verified that $(f,\phi)$ is $\frac{L_{f}}{\mu_{\phi}}$-smooth adaptable restricted on $\mathcal{X}$, but $(f,\phi)$ may not be $L$-smooth adaptable on $\mathrm{int}\,\mathrm{dom}\,\phi$ according to the original definition in [10, Section 2.2] because $\phi$ may not be strongly convex on its whole interior. For example, the entropy function $\phi(x)=\sum^{n}_{i=1}x_{i}(\log x_{i}-1)$ is $\frac{1}{\alpha}$-strongly convex on $[0,\,\alpha]^{n}$ with any $\alpha>0$, but it is not strongly convex on $\mathrm{int}\,\mathrm{dom}\,\phi=\mathbb{R}^{n}_{++}$. Therefore, adopting the notion of restricted $L$-smooth adaptable property in Definition 2.1 could broaden the potential applications of the algorithm developed in this work.
@@ 3 An inexact Bregman proximal difference-of-convex algorithm
Note from the iBPDCA in Algorithm 1 that, at each iteration, our inexact framework allows one to approximately solve the subproblem (3.1) under condition (3.2) satisfying one of the relative stopping criteria (SC1) and (SC2). Since the subproblem (3.1) admits a solution $\boldsymbol{x}^{k,*}\in\mathrm{dom}\,P_{1}\,\cap\,\mathrm{int}\,\mathrm{dom}\,\phi$ (by Assumption A5), then condition (3.2) always holds at $\boldsymbol{x}^{k+1}=\boldsymbol{x}^{k,*}$ with $\|\Delta^{k}\|=\delta_{k}=0$ and hence it is achievable. Moreover, when setting $\sigma=0$ in (SC1) or (SC2), it means that $\boldsymbol{x}^{k+1}$ is an exact solution of the subproblem, and thus, our iBPDCA readily reduces to an exact BPDCA studied in [42].
@@
Combining above two inequalities, we have that
@@
    The sequence $\{F(\bm{x}^{k})\}_{k=0}^{\infty}$ is non-increasing and $\zeta:=\lim\limits_{n\to\infty}F(\bm{x}^{k})$ exists.
@@
    The sequence $\big{\{}F(\bm{x}^{k})+\sigma\gamma_{\max}\mathcal{D}_{\phi}(\bm{x}^{k},\,\bm{x}^{k-1})\big{\}}_{k=0}^{\infty}$ is non-increasing, and $\widetilde{\zeta}:=\lim\limits_{n\to\infty}F(\bm{x}^{k})+\sigma\gamma_{\max}\mathcal{D}_{\phi}(\bm{x}^{k},\,\bm{x}^{k-1})$ exists.

### v2

@@ Abstract
In this paper, we consider a class of difference-of-convex (DC) optimization problems, which require only a weaker restricted $L$-smooth adaptable property on the smooth part of the objective function, instead of the standard global Lipschitz gradient continuity assumption. Such problems are prevalent in many contemporary applications such as compressed sensing, statistical regression, and machine learning, and can be solved by a general Bregman proximal DC algorithm (BPDCA). However, the existing BPDCA is developed based on the stringent requirement that the involved subproblems must be solved exactly, which is often impractical and limits the applicability of the BPDCA. To facilitate the practical implementations and wider applications of the BPDCA, we develop an inexact Bregman proximal difference-of-convex algorithm (iBPDCA) by incorporating two types of relative-type stopping criteria for solving the subproblems. The proposed inexact framework has considerable flexibility to encompass many existing exact and inexact methods, and can accommodate different types of errors that may occur when solving the subproblem. This enables the potential application of our inexact framework across different DC decompositions to facilitate the design of a more efficient DCA scheme in practice. The global subsequential convergence and the global sequential convergence of our iBPDCA are established under suitable conditions including the Kurdyka-Łojasiewicz property. Some numerical experiments on the $\ell_{1-2}$ regularized least squares problem and the constrained $\ell_{1-2}$ sparse optimization problem are conducted to show the superior performance of our iBPDCA in comparison to existing algorithms. These results also empirically validate the necessity and significance of developing different types of stopping criteria to facilitate the efficient computation of the subproblem in each iteration of our iBPDCA.
@@ 1 Introduction
Due to the significance and popularity of the problem (1.1), tremendous efforts have been made to solve it efficiently, particularly when the problem involves a large number of variables. A classical algorithm for solving problem (1.1) is the so-called DC algorithm (DCA) [37], which iteratively replaces the concave part of the objective with a linear majorant approximation and solves the resulting subproblem as follows: $\bm{x}^{k+1}\in\min\limits_{\bm{x}\in{\mathcal{Q}}}\left\{P_{1}(\bm{x})+f(\bm{x})-\langle\bm{\xi}^{k},\,\bm{x}-\bm{x}^{k}\rangle\right\},$ where $\bm{\xi}^{k}\in\partial P_{2}(\bm{x}^{k})$ is a (classical) subgradient of $P_{2}$ at $\bm{x}^{k}\in\mathcal{Q}$ (see the next section for the definition). But solving this subproblem is generally computationally demanding. By further exploring the structures of the objective, a so-called proximal DC algorithm (pDCA) (see, e.g, [17, 38]) can be adopted to solve problem (1.1), which, in each iteration, not only replaces the concave part of the objective with a linear majorant approximation, but also replaces the smooth part with a quadratic majorant approximation. The basic iterative step of the pDCA reads as follows:
@@
The success of the pDCA and the pDCAe also motivates many subsequent works focusing on them and their variants for solving problem (1.1) under appropriate settings; see, for example, [29, 30, 36, 39]. Among them, Takahashi et al. [45] considered the Bregman distance $\mathcal{D}_{\phi}$ (defined in the next section) associated with a kernel function $\phi$ as a proximity measure, and generalized the pDCA to a Bregman proximal DC algorithm (BPDCA), whose basic iterative step reads as follows:
@@
where $\bm{\xi}^{k}\in\partial P_{2}\left(\bm{x}^{k}\right)$. The development of the BPDCA is indeed inspired by recent significant advancements in the Bregman proximal gradient method (see, e.g, [5, 10, 28, 48]). In comparison to the pDCA, such an extended Bregman framework presents two notable advantages. First, the BPDCA can be developed based on the notion of $L$-smooth adaptable property [10, Definition 2.2], which is less restrictive than the global Lipschitz gradient continuity of $f$ that is required by the pDCA(e) and its variants. Therefore, the BPDCA is applicable for a wider range of problems. Second, by selecting an appropriate kernel function $\phi$, the BPDCA can better leverage the inherent geometry and structure of the problem, potentially leading to a more tractable subproblem (1.2); see examples in [45, 46].
@@
Despite the aforementioned advantages, the BPDCA still encounters limitations in practical implementations. Indeed, although the kernel function can possibly capture the inherent geometry and structure of the problem, the associated subproblem in form of (1.2) generally has no closed-form solution and its computation may still be numerically demanding; see examples in [12, 42, 52] and our numerical section 5. Therefore, for the BPDCA to be implementable and practical, it should allow approximate solutions to the subproblem with progressive accuracy and the corresponding stopping criterion should be practically verifiable. However, to our knowledge, research on the inexact version of the BPDCA remains unexplored. On the other hand, allowing the approximate minimization of the subproblem in proximal DC-type algorithms using the classical or weighted quadratic proximal term (corresponding to $\phi(\cdot):=\frac{1}{2}\|\cdot\|^{2}$ or $\phi(\cdot):=\frac{1}{2}\|\cdot\|^{2}_{H}$ with a symmetric positive definite matrix $H$) has been widely investigated under different (typically simpler) contexts in previous works. However, these inexact versions either require the exact computation of an element in $\partial P_{1}$ (e.g., [26, 32, 43, 47, 59]), which could be difficult to satisfy when $P_{1}$ is not a simple function (see an example in the numerical section 5.2), or they lack the global sequential convergence result (e.g., [14, 44, 60]). Moreover, all these inexact versions require the global Lipschitz continuity of $\nabla f$ when they explicitly approximate the smooth part $f$ by a quadratic majorant. The aforementioned inadequacies could potentially restrict the applicability of these existing inexact frameworks, even under the Euclidean setting.
@@
In this paper, to facilitate the practical implementations of the BPDCA and complement the aforementioned existing inexact proximal DC-type algorithms, we attempt to develop a general inexact Bregman proximal DC algorithm (iBPDCA) for solving the general problem (1.1). As we shall see later in Section 3, our inexact framework is developed based on two types of relative stopping criteria (SC1) and (SC2), which vary in their strategy to control the error incurred in the inexact minimization of the subproblem. Although the difference may look subtle at the first glance, it indeed influences the behavior of our iBPDCA in both theoretical analysis and computational performance; see more discussions following Algorithm 1 as well as Remarks 5.1 and 5.2. Moreover, the resulting inexact Bregman proximal DCA framework is rather broad to cover many existing exact and inexact methods (studied in, e.g., [43, 44, 45, 47, 59]), while also introducing new inexact variants. Additionally, it can accommodate different types of errors that may occur when solving the subproblem. The inherent versatility and flexibility of our iBPDCA readily enable its potential application across different DC decompositions for a given problem, thereby facilitating the design of a more efficient DCA scheme, as deliberated in Remark 5.1. Furthermore, in Section 4, we meticulously establish both global subsequential and global sequential convergence results for our iBPDCA under suitable conditions including the renown Kurdyka-Łojasiewicz property. Finally, we conduct some numerical experiments in Section 5 to illustrate the performance of our iBPDCA for solving the $\ell_{1-2}$ regularized least squares problem and the constrained $\ell_{1-2}$ sparse optimization problem. The computational results demonstrate the superior performance of our inexact framework in comparison to existing algorithms, and they also empirically validate the necessity and significance of developing different types of stopping criteria for solving the subproblem.
@@ 2 Notation and preliminaries
In this paper, we present scalars, vectors, and matrices in lower case letters, bold lower case letters, and upper case letters, respectively. We also use $\mathbb{R}$, $\mathbb{R}^{n}$, and $\mathbb{R}^{m\times n}$ to denote the set of real numbers, $n$-dimensional real vectors, and $m\times n$ real matrices, respectively. For a vector $\bm{x}\in\mathbb{R}^{n}$, $x_{i}$ denotes its $i$-th entry, $\|\bm{x}\|$ denotes its Euclidean norm, $\|\bm{x}\|_{1}$ denotes its $\ell_{1}$ norm defined by $\|\bm{x}\|_{1}:=\sum^{n}_{i=1}|x_{i}|$, and $\|\bm{x}\|_{H}:=\sqrt{\langle\bm{x},\,H\bm{x}\rangle}$ denotes its weighted norm associated with a symmetric positive definite matrix $H$. For a closed set $\mathcal{X}\subseteq\mathbb{E}$, its indicator function $\iota_{\mathcal{X}}$ is defined by $\iota_{\mathcal{X}}(\bm{x})=0$ if $\bm{x}\in\mathcal{X}$ and $\iota_{\mathcal{X}}(\bm{x})=+\infty$ otherwise. The distance from a point $\bm{x}$ to $\mathcal{X}$ is defined by $\mathrm{dist}(\bm{x},\,\mathcal{X}):=\inf_{\bm{y}\in\mathcal{X}}\|\bm{y}-\bm{x}\|$.
@@
When $h$ is continuously differentiable or convex, the above subdifferential coincides with the classical concept of derivative or convex subdifferential of $h$; see, for example, [41, Exercise 8.8] and [41, Proposition 8.12]. Moreover, for a proper closed convex function $h:\mathbb{E}\rightarrow(-\infty,\infty]$ and a given $\varepsilon\geq 0$, the $\varepsilon$-subdifferential of $h$ at $\bm{x}\in{\rm dom}\,h$ is defined by $\partial_{\varepsilon}h(\bm{x}):=\{\bm{d}\in\mathbb{E}:h(\bm{y})\geq h(\bm{x})+\langle\bm{d},\,\bm{y}-\bm{x}\rangle-\varepsilon,~{}\forall\,\bm{y}\in\mathbb{E}\}$, and when $\varepsilon=0$, $\partial_{\varepsilon}h$ simply reduces to the classical convex subdifferential $\partial h$.
The function $h$ is essentially smooth if (i) $\mathrm{int}\,\mathrm{dom}\,h$ is non-empty; (ii) $h$ is differentiable on $\mathrm{int}\,\mathrm{dom}\,h$; (iii) $\|\nabla h(\bm{x}^{k})\|\to\infty$ for every sequence $\{\bm{x}^{k}\}$ in $\mathrm{int}\,\mathrm{dom}\,h$ converging to a boundary point of $\mathrm{int}\,\mathrm{dom}\,h$; see [40, page 251]. For a proper closed function $h:\mathbb{E}\rightarrow(-\infty,\infty]$ and $\nu>0$, the proximal mapping of $\nu h$ at $\bm{y}\in\mathbb{E}$ is defined by
@@
Some remarks are in order concerning this definition. The above restricted $L$-smooth adaptable property is derived by modifying the original $L$-smooth adaptable property introduced in [10, Section 2.2] by imposing a restricted set $\mathcal{X}$, and it readily reduces to the original notion when $\mathcal{X}$ is the entire space $\mathbb{E}$. Implementing such a restriction would facilitate the expansion of the original notion to encompass a broader range of choices of $(f,\,\phi)$ with a proper $\mathcal{X}$. For example, when $\nabla f$ is $L_{f}$-Lipschitz continuous on $\mathbb{R}^{n}$ and $\phi$ is $\mu_{\phi}$-strongly convex on $\mathcal{X}\subseteq\mathbb{R}^{n}$, then it can be verified that $(f,\phi)$ is $\frac{L_{f}}{\mu_{\phi}}$-smooth adaptable restricted on $\mathcal{X}$, but $(f,\phi)$ may not be $L$-smooth adaptable on $\mathrm{int}\,\mathrm{dom}\,\phi$ according to the original definition in [10, Section 2.2] because $\phi$ may not be strongly convex on its whole interior. For example, the entropy function $\phi(x)=\sum^{n}_{i=1}x_{i}(\log x_{i}-1)$ is $\frac{1}{\alpha}$-strongly convex on $[0,\,\alpha]^{n}$ with any $\alpha>0$, but it is not strongly convex on $\mathrm{int}\,\mathrm{dom}\,\phi=\mathbb{R}^{n}_{++}$. Therefore, adopting the notion of restricted $L$-smooth adaptable property in Definition 2.1 could broaden the potential applications of the algorithm developed in this work.
@@ 3 An inexact Bregman proximal difference-of-convex algorithm
Note from the iBPDCA in Algorithm 1 that, at each iteration, our inexact framework allows one to approximately solve the subproblem (3.1) under condition (3.2) satisfying one of the relative stopping criteria (SC1) and (SC2). Since the subproblem (3.1) admits a solution $\bm{x}^{k,*}\in\mathrm{dom}\,P_{1}\,\cap\,\mathrm{int}\,\mathrm{dom}\,\phi$ (by Assumption A5), then condition (3.2) always holds at $\bm{x}^{k+1}=\bm{x}^{k,*}$ with $\|\Delta^{k}\|=\delta_{k}=0$ and hence it is achievable. Moreover, when setting $\sigma=0$ in (SC1) or (SC2), it means that $\bm{x}^{k+1}$ is an exact solution of the subproblem, and thus, our iBPDCA readily reduces to an exact BPDCA studied in [45].
@@
Combining above two relations, we have that
@@
    The sequence $\{F(\bm{x}^{k})\}_{k=0}^{\infty}$ is non-increasing and $\zeta:=\lim\limits_{k\to\infty}F(\bm{x}^{k})$ exists.
@@
    The sequence $\big{\{}F(\bm{x}^{k})+\sigma\gamma_{\max}\mathcal{D}_{\phi}(\bm{x}^{k},\,\bm{x}^{k-1})\big{\}}_{k=0}^{\infty}$ is non-increasing, and $\widetilde{\zeta}:=\lim\limits_{k\to\infty}F(\bm{x}^{k})+\sigma\gamma_{\max}\mathcal{D}_{\phi}(\bm{x}^{k},\,\bm{x}^{k-1})$ exists.